The perimeter of a rectangular farm is 10km.its length is one and half times its breadth.find the length and breadth of the farm
Answer by stanbon(75887)

The perimeter of a rectangular farm is 10km.its length is one and half times its breadth.find the length and breadth of the farm
-------
P = 2(length + width)
-----
10 km = 2((3/2)w + w)
----
5 = (5/2)w
w = 2 km (breadth)
---
length = (3/2)w = 3 km
----------
cheers,
Stan H.
----------